OpenStudy (anonymous):

after 4 apples were added to a sack, there were 3 times as many apples in the sack as before. how many apples were in the sack before the addition?

OpenStudy (anonymous):

There was only one apple in it to begin with.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

2+4 = 6 6/3 = 2 so yea. 2. mental math :D

OpenStudy (radar):

Use algebra. Let x = Original number of apples in sack x + 4 = number of apples after 4 were added. Now write the relationship that is given. x + 4 = 3x 2x = 4 x = 2

OpenStudy (radar):

double check original amount was 2 add 4 more getting 6 and 6 is three times 2 everything is good.

OpenStudy (radar):

@redpanda123 it is good to ask why!

OpenStudy (anonymous):

oh thank you